OpenStudy (anonymous):

ln(2t+7)=ln3-ln(t+1) Please help me solve for x.

OpenStudy (anonymous):

You mean solve for t? Let's walk thru it together. Do you know the properties of logarithms?

OpenStudy (anonymous):

Oops yeah t .. I think I do.

OpenStudy (anonymous):

Ok lets use the quotient property on the right hand side of the equation. ln(2t+7) = ln(3/t+1)

OpenStudy (anonymous):

Do you know how i got this?

OpenStudy (anonymous):

Yes sir

OpenStudy (anonymous):

Now I can just drop the ln. Correct?

OpenStudy (anonymous):

Yes you can I was looking for the name of the property, but can't find it. But yes that what you do drop the 'ln' and solve for 't'

OpenStudy (anonymous):

So you should have: 2t+7 = 3/(t+1). Do you know how to find t from here?

OpenStudy (anonymous):

Yes . Thank you for your help.

OpenStudy (anonymous):

Sure no problem.